# AT2018fyk: Candidate Tidal Disruption Event by a (Super)massive Black Hole Binary

### v1

@@ Appendix A Constraints of $M_{\bullet}$ and $a_{\bullet}$ using different X-ray spectra
The slimd model is based on the work of Wen et al. (2020) and Wen et al. (2021), who used a general relativistic stationary slim disk model (Abramowicz et al., 1996; Sądowski, 2009) employing the ray-tracing code from Psaltis & Johannsen (2012) to calculate the synthetic spectrum as seen by an observer. The model assumptions include: 1. A circular and equatorial inner disk. 2. A zero-torque inner boundary condition, large optical depth, and absence of self-irradiation. 3. Generation of X-rays occur at the finite-height photospheric surface of the disk, presumed to match its scale height. The X-ray spectrum is determined by BH mass and spin, the mass accretion rate, the system inclination, and the spectral hardening factor $f_{c}$. The value of the latter parameter is determined through radiative transfer calculations by Davis & El-Abd (2019). Studies have indicated that different choices of the hardening factor have a limited impact on the estimation of mass and spin (see figures 14 and 15 in Wen et al. 2020). To study the constraining power on disk parameters of the thermal spectrum, we fit the Swift 1 spectrum with the slimd model.

### v3

@@ Appendix A Constraints of $M_{\bullet}$ and $a_{\bullet}$ using different X-ray spectra
The slimd model (Wen et al., 2022) is based on the work of Wen et al. (2020) and Wen et al. (2021), who used a general relativistic stationary slim disk model (Abramowicz et al., 1996; Sądowski, 2009) employing the ray-tracing code from Psaltis & Johannsen (2012) to calculate the synthetic spectrum as seen by an observer. The model assumptions include: 1. A circular and equatorial inner disk. 2. A zero-torque inner boundary condition, large optical depth, and absence of self-irradiation. 3. Generation of X-rays occur at the finite-height photospheric surface of the disk, presumed to match its scale height. The X-ray spectrum is determined by BH mass and spin, the mass accretion rate, the system inclination, and the spectral hardening factor $f_{c}$. The value of the latter parameter is determined through radiative transfer calculations by Davis & El-Abd (2019). Studies have indicated that different choices of the hardening factor have a limited impact on the estimation of mass and spin (see figures 14 and 15 in Wen et al. 2020). To study the constraining power on disk parameters of the thermal spectrum, we fit the Swift 1 spectrum with the slimd model.